OpenStudy (anonymous):

I Just Need The Answer. Determine whether line segment FG is tangent to circle E. True False

OpenStudy (rob1525):

need a pic.

OpenStudy (anonymous):

I'm so sorry! I will put it up right away...I forgot

OpenStudy (anonymous):

if angle FGE is 90% then the answer is yes, if not the answer is no. How do we determine if FGE is 90% given the length of 3 sides?

OpenStudy (anonymous):

I guess we could mesure it wilh something

OpenStudy (anonymous):

Alternatively, let's assume that the angle is 90 making the triangle a right hand triangle. If that is true then A^2 + B^2 = C^2 You are given the lengths A,B and C in the diagram. Do things add up correctly? Does 17^2 + 26^2 = 30^2 ?

OpenStudy (rob1525):

Doesnt' work out does it.